OpenStudy (anonymous):

PLEASE HELP write a linear equation in slope-intercept form to model the situation. Ataxi driver charges 5$ plus $0.30 per mile

jimthompson5910 (jim_thompson5910):

let x = # of miles driven

jimthompson5910 (jim_thompson5910):

if it costs $0.30 per mile, then driving x miles will mean you are charged 0.30x dollars so far. Then you have to add on the fixed cost of $5 since you are charged this no matter how far you go (2 miles or 100 miles) to get 0.30x + 5 This means that if y is the total cost, then y = 0.30x + 5

OpenStudy (anonymous):

the choices are a: m=5+0.3c b:C=0.3+5m c: C=5-0.3m d: C=5+0.3m

jimthompson5910 (jim_thompson5910):

remember that 0.30 = 0.3 and all you have to do is replace x with m and y with c m = # of miles c = total cost

OpenStudy (anonymous):

so B?

jimthompson5910 (jim_thompson5910):

look at y = 0.30x + 5 more closely